# Regularity and temperature of stationary black hole event horizons

### v1

@@ 4 Existence of cross-sections
thus after a cycle the tangent vector gets expanded by $e^{-c}$. If $c<0$ we have future incompleteness and past completeness for all generators, if $c>0$ we have future completeness and past incompleteness for all generators, and if $c=0$ we have completeness in both directions for all generators. If $c\neq 0$ we speak of non-degenerate case and if $c=0$ of degenerate case.
Notice that the formulation of this conclusion on the existence of $c$ is independent of rescalings of $n$ (and hence of the $S^{1}$-generating property).
Let us consider the non-degenerate case. By the results in [45, 20] we know that $n$ can be chosen so that it has constant surface gravity. Actually, in this simplified setting of trivial $S^{1}$-bundle this result can be obtained directly by specializing the formulas obtained in [20]. Let $n$ be a $S^{1}$-generating field with period $\tau$. Recalling that $\hat{H}$ is diffeomorphic to $S^{1}\times\Sigma$, we define the function on $\hat{H}$
@@
where $x_{q}$ is the integral curve of $n$ starting from $q$. The first expression shows that the function is periodic $f(x_{q}(r))=f(x_{q}(r+\tau))$ and so well defined. The second expression shows that the surface gravity of $n^{\prime}=e^{f}n$ is $\kappa^{\prime}=e^{f}(\partial_{n}f+\kappa)=c/\tau$.
So let $n$ be any tangent field with constant non-zero surface gravity. We can still calculate the dilating factor as done previously to get
@@
where this time the range of the parameter might depend on the generator as $n$ is not $S^{1}$-generating a priori. But the just proved equation proves that $R(\gamma)=c/\kappa$, so it is indeed independent of $\gamma$. By rescaling $n$ with a global constant we can accomplish $R=\tau$. We have just proved that, up to a constant, every choice $n$ of constant non-zero surface gravity is automatically $S^{1}$-generating. If we use the choice $n^{\prime}$ constructed in the previous paragraph, this is precisely $S^{1}$-generating with period $\tau$. This is the choice we make in the non-denegerate case.
@@
In the degenerate case the analogous conclusion is easier to obtain. Let us define the field over a smooth section and propagate it with the geodesic property. This extends to a well defined field over $\hat{H}$ since $e^{-c}=1$. The new field has zero surface gravity. But we have still the freedom of choosing the initial field. Let $\lambda(\gamma)$ be the affine length of the first cycle, which depends on the generator chosen. Rescaling $n\to n\lambda(\gamma)/\tau$ we get that the new $n$ has zero surface gravity and is $S^{1}$-generating with period $\tau$ (one could also start from a $S^{1}$-generating field with period $\tau$ and use the previous formulas for $e^{f}$ in the limit $c\to 0$).
The statement that $\kappa$ and all the vectors fields $n$ for any possible choice of $\tau$ coincide when lifted to $H^{+}$ is proved as follows. First all lifted fields $n$ obtained with our construction are complete thus by Theorem 4.11 the sign of $\kappa$ is well defined.
If $\check{\tau}=0$ the Killing field $k$ is proportional to $n$, thus the condition $\varphi_{\tau}=\phi_{\tau}$ implies $k=n$. This shows that $n$ is fixed and the pregeodesic formula $\nabla_{k}k=\kappa k$ gives the value of $\kappa$ which thus follows directly from the initial data given by $k$. Suppose $\check{\tau}>0$, then $\tau=i\check{\tau}$ for some integer $i>0$. The dilating factor is $(e^{\check{c}})^{i}$, that is $c=i\check{c}$.
There are two cases. If $\kappa\neq 0$ the general formula $c=\kappa\tau$ shows that $\kappa$ not only is univocally determined given a choice of $\tau$ but also that it remains the same for all $i$ and so does $n$ once lifted to $H^{+}$ by Theorem 4.11. If $\kappa=0$ the problem of determining its value does not apply, and for a choice of $i$ the field $n$ is chosen so that the parameter length over the cycle on $\hat{H}$ is $\tau=i\check{\tau}$. Note that this construction can be performed with $i=1$ an the lifted $n$ on $H^{+}$ is a geodesic field such that $p$ and $\phi_{\check{\tau}}(p)$ are separated by affine length $\check{\tau}$. Since the dilating factor is 1, it follows that $p$ and $\phi_{i\check{\tau}}(p)$ are separated by affine length $\tau=i\check{\tau}$. This is the property that would be obtained for $n$ if we were to use the compactification for $\tau=i\check{\tau}$, thus the lifted fields $n$ in $H^{+}$ are the same for all $i$.
@@
Since the flow of $k$ on $\hat{H}$ sends generators to generators, $k$ is projectable under the map $\tilde{\pi}:\hat{H}\to\tilde{S}$. But $\tilde{\pi}_{*}(n)=0$, thus $0=[\tilde{\pi}_{*}(k),\tilde{\pi}_{*}(n)]=\tilde{\pi}_{*}([k,n])$ which implies $[k,n]=fn$ for some function $f:\hat{H}\to\mathbb{R}$. But since $L_{k}\nabla=0$ we have, taking the Lie derivative of (1), $\nabla_{fn}n+\nabla_{n}(fn)=\kappa fn$ which implies $\partial_{n}f+\kappa f=0$ and hence along a generator $\gamma(s)$, $f=C\exp(-\kappa s)$ where $s$ is a parameter such that $n={\rm d}/{\rm d}s$. But since the generator closes itself after a period $\check{\tau}$, $f(\gamma(0))=f(\gamma(\check{\tau}))$ which implies $C=C\exp(-\kappa\check{\tau})$. The only solution for $\kappa\neq 0$ is $C=0$, that is $f=0$ over every generator, and hence $f=0$ over $\hat{H}$.
As for the case $\kappa=0$ we proceed as follows (a different proof will be provided in the last paragraph). Recall that the bundle $\tilde{\pi}:\hat{H}\to\tilde{S}$ is trivial, thus we can find a 1-form $\omega$ such that $\omega(n)=1$, $L_{n}\omega=0$ (in the trivialization $\hat{H}\to S^{1}_{\theta}\times\tilde{S}$, let $\omega={\rm d}\theta$ where $n=\frac{\partial}{\partial\theta}$), so introduced a Riemannian metric $\sigma$ on $\tilde{S}$, $n$ is Killing for the Riemannian metric $\hat{g}:=\tilde{\pi}^{*}\sigma+\omega\otimes\omega$ and normalized, $\hat{g}(n,n)=1$. From $[k,n]=fn$, we have
$\hat{g}(\hat{\nabla}_{k}n,n)-\hat{g}(\hat{\nabla}_{n}k,n)=f.$
As $n$ is a unit Killing vector field for $(\hat{H},\hat{g})$, ^$\hat{\nabla}_{n}n=0$, so we get $-n(\hat{g}(k,n))=f$. From $[k,n]=fn$, we also have
$\hat{g}(\hat{\nabla}_{k}n,k)-\hat{g}(\hat{\nabla}_{n}k,k)=f\hat{g}(k,n).$
Since $n$ is a Killing vector field for $(\hat{H},\hat{g})$, we get $n(\hat{g}(k,k))=-2f\hat{g}(k,n).$ From this we get $n(n(\hat{g}(k,k)))=-2fn(\hat{g}(k,n)$ since, recalling a previous equation, $n(f)=-\kappa f=0$. But $-n(\hat{g}(k,n))=f$, so it follows that $n(n(\hat{g}(k,k)))=2f^{2}$. Let us set $h:=\hat{g}(k,k):\hat{H}\to\mathbb{R}$ which, being continuous and defined on a compact set, is a bounded. For any integral curve $\gamma:\mathbb{R}\to\hat{H}$ of $n$, we have $(h\circ\gamma)^{\prime\prime}(t)\geq 0,\forall\,t\in\mathbb{R};$ hence $h\circ\gamma$ is a convex function defined on $\mathbb{R}$, but since $h\circ\gamma$ is bounded, it is constant. It follows that for any integral curve $\gamma$ of $n$, $(h\circ\gamma)^{\prime\prime}(t)=0,\forall\,t\in\mathbb{R}$; which proves that $f=0$. Hence $[k,n]=0$.
@@
The statement on the axisymmetry follows from the formula for the flow of a linear combination of vector fields that commute $\Phi_{\frac{\check{\tau}}{2\pi}(k-n)t}=\Phi_{\frac{\check{\tau}}{2\pi}kt}\circ\Phi_{-\frac{\check{\tau}}{2\pi}nt}=\phi_{\frac{\check{\tau}}{2\pi}t}\circ\varphi^{-1}_{\frac{\check{\tau}}{2\pi}t}$ which for $t=2\pi$ gives the identity.
@@
Since $k$ is Killing and $H^{+}$ is totally geodesic, it is clear that its flow preserves the induced metric. It is also well known that the flow of $n$ preserves the induced metric (e.g. [16, Lemma B1][38][20, Lemma 7]). Thus the flow of $\zeta$ preserves the induced metric. This means that over every non-trivial orbit (which is closed) $\zeta$ has the same square $g(\zeta,\zeta)$ at each of its points and hence that $\zeta$ has the same causal character at every point of the orbit. It cannot be lightlike (it cannot be timelike as it is tangent to the horizon) as there would be a closed causal curve in contradiction with the strong casuality of $H^{+}$, thus $\zeta$ is spacelike wherever it is non-zero.
@@
For the last statement, we already know by Prop. 4.7 that the bundle $\tilde{\pi}:H^{+}\to\tilde{S}$ is trivial and admits a smooth section. We can thus find a sujective function $f:H^{+}\to\mathbb{R}$ (essentially the first projection in the trivialization) such that ${\rm d}f(n)=1$, and the level sets of $f$ are smooth sections for the bundle. Let $\psi_{s}$ be the flow of $\zeta$. As $\psi$ commutes with $\varphi$ it sends (parametrized) generators to (parametrized) generators, so that $(\psi_{s})_{*}(n)=n$. In particular, $f\circ\psi_{s}$ increases over the generators.
@@
It is smooth and such that ${\rm d}\bar{f}(n)=1$ which, by the completeness of $n$ implies that it is surjective. Furthermore, due to reparametrization invariance $\psi_{\theta}^{*}\bar{f}=\bar{f}$, we have ${\rm d}\bar{f}(\zeta)=0$, which implies that the levels sets of $\bar{f}$ intersect the generators exactly once and are tangent to the orbits of the $S^{1}$ action $\psi$. Each of the level sets is a cross-section.
By construction the cross-section is transverse to $n$ (and so $k$, wherever $k$ is proportional to $n$). But it is also transverse to $k$ elsewhere because $k=n+\frac{2\pi}{\tau}\zeta$ as $\zeta$ is tangent to the surface by construction. The fact that the level sets are preserved by the $\varphi$ flow follows from ${\rm d}\bar{f}(n)=1$. The fact that they are preserved by the $\phi$ flow follows from the fact that any map $\phi_{s}$ can be written as a composition of maps $\phi$ and $\psi$, where the level sets are sent to level sets by both types of maps. ∎
@@
If the black hole is rotating then we can read the theorem in the ‘check’ version, with $\tau=\check{\tau}>0$. The notion of cycle of the geodesic in $\hat{H}$ is entirely geometrical as does the scale factor $e^{\check{c}}$. The modulus of $k$ is supposedly fixed: for instance if the spacetime approximates Minkowski at infinity one typically demands that $k$ should represent the velocity field of an observer in Minkowski, i.e. it should approach a normalized field $g(k,k)\to-1$. This shows that, physically speaking, $k$ cannot be rescaled. It follows that $\check{\tau}$ and $\check{\omega}$ are completely fixed by the geometry. As shown by the theorem, $\kappa$ and the field $n$ are also uniquely determined.
@@
This important geometrical fact confirms the possibility of interpreting surface gravity as a physical quantity, i.e. the temperature of the black hole $T=\frac{\kappa}{2\pi}$.
@@
Thus the previous theorem clarifies that the geometry fixes the value of surface gravity not just its sign. If surface gravity were introduced via the property (1) or with (1) and (2) as done e.g. in [41, Def. 1.22] only the sign of $\kappa$ would be definable as any point independent rescaling of $n$ would be allowed but would change its modulus. It is the condition $\phi_{\tau}=\varphi_{\tau}$, and hence the request that $n$ should be $S^{1}$-generating, that plays a major role in determining the value of $\kappa$.
@@
Notice that we are able to prove that stationarity implies the existence of an $S^{1}$ action on the horizon, which we rightfully term axisymmetry, without extending the involved fields to Killing fields (in fact our proof is different from previous proofs). It seems remarkable that we can obtain the interesting physical quantities of (constant) temperature and angular velocity directly from the geometry of the horizon without using the Einstein’s equations or any extension to Killing fields (a result which requires those equations). Our proof of axisymmetry is also largely independent of dimensionality assumptions.
@@
As mentioned in the above theorem, under the dominant energy condition the totally geodesic compact smooth horizon $\hat{H}$ can either be degenerate ($\kappa=0$) or non-degenerate ($\kappa\neq 0$), namely admit a complete generator or not. This can be seen as a special case of the dichotomy proved in [45, 20].
@@
Further, by recent results by Petersen and Rácz [43, 42, 41] in the vacuum non-degenerete case the tangent field $n$ that realizes the constant surface gravity can be extended to a Killing vector field (possibly different from $k$). By lifting it to the neighborhood $V$ of $H^{+}$ one gets another proof of Hawking’s local rigidity theorem in the non analytic case, which was recently proven by Alexakis-Ionescu-Klainerman [1] using a different analysis taking advantage of a bifurcate horizon assumption. The mentioned compactification strategy has instead been followed by Petersen [43] (who relied on Chruściel-Costa [7] for regularity, so he assumed existence of cross-sections).
@@
We note that in 4 spacetime dimension and for horizons having a section of non-zero Euler characteristic, due to Thm. 4.14, there is no need to work in the most general non-closed generators case as done in [43]. Our result proves the condition $[k,n]=0$ on $H^{+}$ that Petersen had to assume in his work and which was not established in [45, 20]. Thanks to this proof we fill the gap between results on the constancy of surface gravity (that do not mention $[k,n]=0$) and the assumption in Petersen’s work.

### v2

@@ 7 The surface gravity (temperature) and angular velocity of event horizons
thus after a cycle the tangent vector gets expanded by $e^{-c}$. If $c<0$ we have future incompleteness and past completeness for all generators (hawking73, , Comment after Prop. 6.4.3), if $c>0$ we have future completeness and past incompleteness for all generators, and if $c=0$ we have completeness in both directions for all generators. If $c\neq 0$ we speak of non-degenerate case and if $c=0$ of degenerate case. The constant $c$ is denoted $\check{c}$ when we choose $\tau=\check{\tau}$, see Thm. 7.1 for the definition of $\check{\tau}$.
Notice that this property of $c$ is formulated using geodesics and so does not mention the $S^{1}$-generating property of $n$. We just used the closure of the generators on $\hat{H}$, but no special parametrization.
Once $n$ is lifted to $H$ the just proved property becomes that expressed as the vector ratio $\dot{\gamma}(\check{\tau})/d\phi_{\check{\tau}}(\dot{\gamma}(0))=e^{-\check{c}}$ in the statement of the theorem, where the choice $\tau=\check{\tau}$ has been made, indeed, note that $\dot{\gamma}(0)=bn(\gamma(0))$ thus $d\phi_{\check{\tau}}(\dot{\gamma}(0))=d\varphi_{\check{\tau}}(\dot{\gamma}(0))=bd\varphi_{\check{\tau}}(n)=bn(\gamma(\check{\tau}))$ and so the ratio in the statement of the theorem is precisely that we calculated above.
Existence of $\bm{S^{1}}$-generating fields on $\bm{\hat{H}}$ of constant surface gravity. Let us consider the non-degenerate case, $c\neq 0$, namely suppose that there is a (past or future) incomplete generator on $\hat{H}$. By the results in reiris21 ; minguzzi21 we know that $n$ can be chosen so that it has non-zero constant surface gravity. Actually, in this simplified setting of trivial $S^{1}$-bundle this result can be obtained directly by specializing the formulas obtained in minguzzi21 . Let $n$ be a $S^{1}$-generating field with period $\tau$. Recalling that $\hat{H}$ is diffeomorphic to $S^{1}\times\Sigma$, we define the function on $\hat{H}$
@@
where $x_{q}$ is the integral curve of $n$ starting from $q\in\Sigma$. The first expression shows that the function is periodic $f(x_{q}(r))=f(x_{q}(r+\tau))$ and so well defined. The second expression shows that the surface gravity of $n^{\prime}=e^{f}n$ is $\kappa^{\prime}=e^{f}(\partial_{n}f+\kappa)=c/\tau$ which is a non-zero constant.
So let $n$ be any tangent field with constant non-zero surface gravity on $\hat{H}$. We can still calculate the dilating factor as done previously to get
@@
where this time the range of the parameter might depend on the generator as $n$ is not $S^{1}$-generating a priori. But the just proved equation proves that $R(\gamma)=c/\kappa$, so it is indeed independent of $\gamma$. By rescaling $n$ with a global constant we can accomplish $R=\tau$.
We have just proved that, in the non-degenerate case, $c\neq 0$, there is a choice of $n$ on $\hat{H}$ of constant non-zero surface gravity and for any such choice $n$ is automatically $S^{1}$-generating.
@@
In the degenerate case, $c=0$, it is similarly possible to construct a $S^{1}$-generating field of zero surface gravity. Let us define the field over a smooth section and propagate it with the geodesic property. This extends to a well defined field over $\hat{H}$ since $e^{-c}=1$. The new field has zero surface gravity. But we have still the freedom of choosing the initial field. Let $\lambda(\gamma)$ be the affine length of the first cycle, which depends on the generator chosen. Rescaling $n\to n\lambda(\gamma)/\tau$ over each generator $\gamma$ we get that the new $n$ has zero surface gravity and is $S^{1}$-generating with period $\tau$ (one could also start from a $S^{1}$-generating field with period $\tau$ and use the previous formulas for $e^{f}$ in the limit $c\to 0$).
The field $n$ with the required properties (a)-(c) is obtained lifting the analogous field from $\hat{H}$. For $\check{\tau}=0$ we just consider the construction of $n$ on $\hat{H}$ of the previous paragraphs choosing any $\tau>0$ for the compactification. For $\check{\tau}>0$ we just consider the construction of $n$ on $\hat{H}$ of the previous paragraphs choosing $\tau=\check{\tau}$ for the compactification.
Uniqueness of $\bm{n}$ and $\bm{\kappa}$ on $\bm{H^{+}}$.
Assume that the future-directed lightlike field $n$ satisfies (a)-(c).
Suppose $\check{\tau}=0$. We already know from Thm. 7.1 that $n=\alpha k$ for some function $\alpha$ which means that the horizon is Killing, and since both fields do not vanish on $H^{+}$, using Thm. 4.1, $\alpha>0$. Since the dominant energy condition holds we know that $\nabla_{k}k=\kappa^{\prime}k$ for some constant $\kappa^{\prime}$, see (chrusciel20, , Thm. 4.3.12). Thus both $n$ and $k$ have constant surface gravity. But the sign of surface gravity must be the same as it is related to the affine completeness of the generators. If the surface gravity does not vanish, as both $n$ and $k$ are complete on $H^{+}$, Theorem 6.1 implies that they differ by a constant. Condition (c) now implies $k=n$. If the surface gravity does vanish then $0=\alpha^{-1}\nabla_{n}n=\nabla_{k}n=(\partial_{k}\alpha)k+\alpha\nabla_{k}k=(\partial_{k}\alpha)k$ which implies that $\alpha$ is a constant over every generator. But this constant must again be 1 due to condition (c). We conclude that for $\check{\tau}=0$, $n=k$, thus $n$ is unique.
Suppose $\check{\tau}>0$. Property (c) implies that all geodesics return to themselves under $\phi_{\tau}$, thus $\tau=i\check{\tau}$ by Thm. 7.1, where $i$ is a positive integer. Let $n^{\prime}$ be a $S^{1}$-generating field of constant surface gravity which we know to exist on $\hat{H}_{\check{\tau}}$ (obvious meaning of the notation). Its lift to $H^{+}$ is denoted in the same way. Both $n$ and $n^{\prime}$ on $H^{+}$ are complete and have constant surface gravity. The signs of the surface gravities actually coincide because they are related to the affine incompleteness of the generators. If the surface gravities do not vanish then Theorem 6.1 proves that $n$ and $n^{\prime}$ differ by a constant factor. As both satisfy $\varphi_{\tau}=\phi_{\tau}$ this factor is unity. If the surface gravity vanish then, writing $n=\alpha n^{\prime}$, $\alpha>0$, $0=\alpha^{-1}\nabla_{n}n=\nabla_{n^{\prime}}n=(\partial_{n^{\prime}}\alpha)n^{\prime}+\alpha\nabla_{n^{\prime}}n^{\prime}=(\partial_{n^{\prime}}\alpha)n^{\prime}$ which implies that $\alpha$ is a constant over every generator. But this constant must again be 1 due to condition $\varphi_{\tau}=\phi_{\tau}$, which proves $n=n^{\prime}$ and so all possible $n$ coincide. In particular, the identity $n=n^{\prime}$ implies $\varphi_{\check{\tau}}=\phi_{\check{\tau}}$.
@@
Commutativity $\bm{[k,n]=0}$. Since the flow of $k$ on $\hat{H}$ sends generators to generators, $k$ is projectable under the map $\tilde{\pi}:\hat{H}\to\tilde{S}$. But $\tilde{\pi}_{*}(n)=0$, thus $0=[\tilde{\pi}_{*}(k),\tilde{\pi}_{*}(n)]=\tilde{\pi}_{*}([k,n])$ which implies $[k,n]=fn$ for some function $f:\hat{H}\to\mathbb{R}$. But since $L_{k}\nabla=0$ we have, taking the Lie derivative of $\nabla_{n}n=\kappa n$, $\nabla_{fn}n+\nabla_{n}(fn)=\kappa fn$ which implies $\partial_{n}f+\kappa f=0$ and hence along a generator $\gamma(s)$, $f=C\exp(-\kappa s)$ where $s$ is a parameter such that $n={\rm d}/{\rm d}s$. But since the generator closes itself after a period $\check{\tau}$, $f(\gamma(0))=f(\gamma(\check{\tau}))$ which implies $C=C\exp(-\kappa\check{\tau})$.
Suppose $\kappa\neq 0$, then the only solution is $C=0$, that is $f=0$ over every generator, and hence $f=0$ over $\hat{H}$.
[…]
@@
The $\bm{S^{1}}$-action on $\bm{H^{+}}$. The statement on the axisymmetry follows from the formula for the flow of a linear combination of vector fields that commute $\Phi_{\frac{\check{\tau}}{2\pi}(k-n)t}=\Phi_{\frac{\check{\tau}}{2\pi}kt}\circ\Phi_{-\frac{\check{\tau}}{2\pi}nt}=\phi_{\frac{\check{\tau}}{2\pi}t}\circ\varphi^{-1}_{\frac{\check{\tau}}{2\pi}t}$ which for $t=2\pi$ gives the identity.
@@
Since $k$ is Killing and $H^{+}$ is totally geodesic, it is clear that its flow preserves the induced metric. It is also well known that the flow of $n$ preserves the induced metric (e.g. (friedrich99, , Lemma B1)moncrief08 (minguzzi21, , Lemma 7)). Thus the flow of $\zeta$ preserves the induced metric. This means that over every non-trivial orbit (which is closed) $\zeta$ has the same square $g(\zeta,\zeta)$ at each of its points and hence that $\zeta$ has the same causal character at every point of the orbit. It cannot be lightlike as there would be a closed causal curve in contradiction with the strong casuality of $H^{+}$, and it cannot be timelike as it is tangent to the horizon, thus $\zeta$ is spacelike wherever it is non-zero.
Existence of nice cross-sections. For the last statement, we already know by Thm. 7.1 that the bundle $\tilde{\pi}:H^{+}\to\tilde{S}$ is trivial and admits a smooth section.
@@
We can thus find a surjective function $f:H^{+}\to\mathbb{R}$ (essentially the first projection in the trivialization) such that ${\rm d}f(n)=1$, and the level sets of $f$ are images of smooth sections for the bundle. If $\check{\tau}=0$ we have $n=k$, $\zeta=0$, and we can set $\bar{f}:=f$. For $\check{\tau}>0$, let $\psi_{s}$ be the flow of $\zeta$. As $\psi$ commutes with $\varphi$ it sends (parametrized) generators to (parametrized) generators, so that $(\psi_{s})_{*}(n)=n$. In particular, for any fixed $s$, $f\circ\psi_{s}$ increases over the generators.
@@
It is smooth and such that ${\rm d}\bar{f}(n)=1$ which, by the completeness of $n$ on $H^{+}$, implies that it is surjective over every generator. This implies that the levels sets of $\bar{f}$ intersect the generators exactly once. Each of the level sets is a cross-section. Since $\phi_{\tau}=\varphi_{\tau}$, $f(\phi_{m\tau})=f(\varphi_{m\tau})$ for every integer $m$ which, by continuity, shows that $f$ is surjective over every $k$-Killing orbit. Furthermore, due to reparametrization invariance $\psi_{\theta}^{*}\bar{f}=\bar{f}$, we have ${\rm d}\bar{f}(\zeta)=0$, which shows that the orbits of the $S^{1}$ action $\psi$ run over the level sets of $\bar{f}$, and due to the expression for $\zeta$, $k=n+\frac{2\pi}{\tau}\zeta$, we have ${\rm d}\bar{f}(k)=1$, which proves that $\bar{f}$ is increasing over every $k$-Killing orbit, and so that the levels sets of $\bar{f}$ intersect the $k$-Killing orbits exactly once.
The fact that the level sets are preserved by the $\varphi$ flow follows from ${\rm d}\bar{f}(n)=1$. The fact that they are preserved by the $\phi$ flow follows from ${\rm d}\bar{f}(k)=1$. $\square$
@@
The Killing property of a field is preserved by multiplication by a global constant. Still, we cannot assign this freedom to the Killing field $k$: for instance, if the spacetime approximates Minkowski at infinity, one typically demands that $k$ should represent the velocity field of an observer in Minkowski, i.e. it should approach a normalized field $g(k,k)\to-1$. This shows that, physically speaking, $k$ cannot be rescaled. It follows that, as show in Theorem 7.1, $\check{\tau}$ and $\check{\omega}$ are completely fixed by the geometry. As shown by Theorem 7.2, $\kappa$ and the field $n$ are also uniquely determined.
@@
This important geometrical fact confirms the possibility of interpreting the surface gravity as a physical quantity, i.e. the temperature of the black hole $T=\frac{\kappa}{2\pi}$.
@@
Thus the previous theorem clarifies that the geometry fixes the value of surface gravity not just its sign. If surface gravity were introduced via the property (3) or with (3) and (4) as done e.g. in (petersen19, , Def. 1.22) only the sign of $\kappa$ would be definable as any point independent rescaling of $n$ would be allowed. It is the condition $\phi_{\tau}=\varphi_{\tau}$, and hence the request that $n$ should be $S^{1}$-generating, that plays a major role in determining the value of $\kappa$.
@@
Notice that we are able to prove that stationarity implies the existence of an $S^{1}$ action on the horizon, which we rightfully term axisymmetry, without extending the involved fields to Killing fields (in fact our proof is different from previous proofs). It seems remarkable that we can obtain the interesting physical quantities of (constant) temperature and angular velocity directly from the geometry of the horizon without using a vacuum assumption or any extension to Killing fields (a result which requires the Einstein’s equations). Our proof of axisymmetry is also largely independent of dimensionality assumptions.
@@
As mentioned in the above theorem, under the dominant energy condition the totally geodesic compact smooth horizon $\hat{H}$ can either be degenerate ($\kappa=0$) or non-degenerate ($\kappa\neq 0$), namely admit a complete generator or not. This can be seen as a special case of the dichotomy proved in reiris21 ; minguzzi21 .
@@
Further, by recent results by Petersen and Rácz petersen18 ; petersen18b ; petersen19 in the vacuum non-degenerate case, the tangent field $n$ that realizes the constant surface gravity can be extended to a Killing vector field (possibly different from $k$). By lifting it to the neighborhood $V$ of $H^{+}$ one gets another proof of Hawking’s local rigidity theorem in the non analytic case, which was recently proven by Alexakis-Ionescu-Klainerman alexakis10 using a different analysis taking advantage of a bifurcate horizon assumption. The mentioned compactification strategy has instead been followed by Petersen petersen18 (who relied on Chruściel-Costa chrusciel08 for regularity, so he assumed existence of cross-sections).
@@
We note that in 4 spacetime dimension and for horizons having a section of non-zero Euler characteristic, due to Thms. 7.1-7.2, there is no need to work in the most general non-closed generators case as done in petersen18 . Our result proves the condition $[k,n]=0$ on $H^{+}$ that Petersen (petersen19, , Thm. 1.23) had to assume in his work and which was not established in reiris21 ; minguzzi21 . Thanks to this proof we fill the gap between previous results on the constancy of surface gravity (that do not mention $[k,n]=0$) and the assumption in Petersen’s and Petersen and Racz’s work.